OpenStudy (anonymous):

A cargo flew to Melbourne and back. It took one hour longer to go there than it did to come back. The average speed on the trip there was 210km/h. The average speed on the way back was 240km/h. How many hours did the trip there take?

OpenStudy (anonymous):

total time for to & fro journey is 7hrs........rgt?

OpenStudy (anonymous):

Hey genuine is it crt?

OpenStudy (ar43r):

let x be time for the trip back 240x=210(x+1) 240x=210x+210 30x=210 x=7 the way there took 1 h more. so 8h